# Tuning-free Visual Effect Transfer across Videos

### v3

@@ 4.2 Baselines
Image-to-Video Baselines. We benchmark against the Wan 2.1 image-to-video model [wan2025wan] and the Wan VACE model [jiang2025vace], which represent current state-of-the-art text-guided video generation systems.
Video-to-Video Baselines. For the video-to-video task, we evaluate two configurations based upon the Wan VACE [jiang2025vace] conditional model and one native text-based video editing model. Pose-conditioned Wan VACE: In this setup, we condition the Wan VACE on the first frame of the input video, intermediate poses extracted from the input video [yang2023effective], and the corresponding text prompt. Depth-conditioned: In this setup, we condition the Wan VACE on the first frame of the input video, intermediate depth maps [ranftl2020towards] from the input video, and the same text prompt. Lucy Edit [Hleihil2025DecartAI]: Lucy Edit is a text-based video editing model built on Wan 2.2 [Hleihil2025DecartAI], which directly takes an input video and a text instruction to produce an edited output.
@@ 4.3 Qualitative Results
Video-to-Video Effects. Qualitative comparisons for the video-to-video task are shown in Figure 7. Wan VACE variants either overfit to the input video or introduce artifacts, while Lucy Edit produces static, frame-invariant edits that lack temporal evolution. By conditioning on the full reference video, our method consistently reproduces complex time-varying transformations—such as gradual color shifts or structural morphing—while maintaining spatial consistency and motion alignment with the input.
@@
Figure 7: Results for our reference-based video-to-video setup. We compare three baselines: (1) Wan VACE + Pose: a conditional diffusion transformer [jiang2025vace] using the first input frame and extracted poses; (2) Wan VACE + Depth: using the first frame and extracted depths with Wan VACE; and (3) Lucy Edit [Hleihil2025DecartAI], a video editing model fine-tuned from an image-to-video model. Without reference video information, all baselines fail to follow the reference from a prompt alone. Wan VACE methods overfit to the input or add irrelevant edits, while Lucy Edit applies uniform static edits across frames. Our method successfully follows the reference effect.
@@
User Study For a subjective task such as reference based video editing or image to video, it is hard to determine any one quantitative metric that fully encompasses success. As a result, we conduct a comprehensive user study as the main way to evaluate our method across three tasks: neural video-to-video effect transfer (V2V), image-to-video effect transfer (I2V), and code-based temporal effects. Annotators performed pairwise comparisons between our method and several baselines, including Lucy Edit, VACE (Depth/Pose/I2V variants), Wan2.1 , and ablations (Use our trained model, but do drop the reference effect video conditioning) , rating each pair on Reference Video Adherence and Input Video Adherence dimensions using a granular 5-level scale. We collected 669 valid annotations for V2V, 455 for I2V, and 252 for code-based effects. Our method demonstrates strong performance in reference adherence, achieving win rates of 78-94% against competitive baselines in V2V, 94-98% in code-based effects, and 75-82% in I2V, as shown in Tables 2 and 1. For input video adherence in V2V, we achieve win rates of 70-98% against method-based baselines (Lucy Edit, VACE Depth/Pose), as shown in Table 1, while comparisons with our ablations show expected behavior: ties with No Style (which uses identical input processing). Mean scores across all dimensions are consistently positive (+0.15 to +1.51), indicating systematic preference for our method’s outputs, and we provide full results in the supplement. These results validate our approach’s effectiveness in effect transfer while preserving input content fidelity across multiple generation modalities. For more details on the user study setup, please refer to the Supplement.
@@ 4.4 Quantitative Metrics
Conditioning Input Similarity. Since our validation sets do not include ground-truth triplets, we measure the similarity of our generated outputs to both the input and reference videos as a proxy for task success. To quantify this, we employ VideoPrism [zhao2024videoprism], a large-scale video embedding model pretrained on over 500M clips, to compute feature-space similarities between videos. For the reference + image-to-video case, we instead measure the average similarity between the generated video and its first-frame conditioning using CLIP [radford2021learning] embeddings. Results are summarized in Tables 3 and 4. Our method consistently achieves higher similarity to the reference video than all baselines, confirming that it effectively incorporates temporal and stylistic information from the reference. Interestingly, baseline models often exhibit slightly higher similarity to the input, likely reflecting a form of under-editing. This phenomenon is also visible in Figure 7 (left) and Figure 6 (bottom), where baselines preserve input structure but fail to reproduce the desired temporal evolution.
@@ 4.5 Controllable Real Video Editing
When editing real videos, users may want to control the level to which the original video is maintained, as well as the amount that the reference effect video effects the final generation. To this end, we follow previous work [ho2022classifier, brooks2023instructpix2pix, jones2024customizing, kumari2025generating, kumari2024customizing, song2025history] and combine multiple classifier-free guidance directions during inference. Specifically, we consider applying guidance for text, input video, and reference effect video, normalizing latents as recomended in previous work [kumari2025generating]. This gives:
@@
Table 3: I2V Similarities. First Frame Sim is average clip [radford2021learning] similarity to the input frame, and Ref Sim. is video embedding similarity [zhao2024videoprism] between the output video and the reference effect video
@@
Table 4: Neural and Code-Based V2V Similarities. Input Sim. is video embedding similarity [zhao2024videoprism] between the output video and the input video, and Ref Sim. is the video embedding similarity between the output video and the reference effect video.
@@ Appendix B Dataset Details
We source the Image-to-Video data from open-source LoRA effects trained on top of the Wan 2.1 Image-to-Video model [wan2025wan], available on Hugging Face [WanI2VLoRAs]. A comprehensive list of LoRA effects and their corresponding captions is provided in Table 5.
We use Qwen Image [wu2025qwen] and Qwen Image Edit [wu2025qwen], together with Wan 2.1 VACE-1.3B [wan2025wan], to implement our text-guided video editing pair creation pipeline, as described in Figure 4 of the main text. We categorize our effect types and general effect categories in Table 6.
For the code-based effects, we source data from the grounding subset of the Senorita dataset [zi2025se]. Specifically, we filter for human-centric videos and include only grounding results with human-centric masks, yielding a collection of over 100K videos. The set of objects considered is listed in Table 7. In the original dataset, input videos are used to predict grounding masks; we repurpose these masks to obtain foreground or object and background segmentations, which we use to augment our effect pipelines. We manually curate code across a diverse set of effects and temporal transitions to sample from when creating triplets of the form (reference video, input video, output video). The full set of effect types and their corresponding hyperparameters is provided in Table 8, and the full set of temporal effect types and their corresponding hyperparameters is given in Table 9.
Appendix C More Implementation Details
@@
Appendix D More Quantitative Results
We further assess video fidelity and perceptual quality using established metrics from VBench [huang2024vbench], including motion smoothness [wan2025wan], aesthetic score [LAION-AI2022aesthetic-predictor], and dynamic degree [teed2020raft, huang2024vbench].
Motion quality measures temporal consistency by computing the average cosine similarity [radford2021learning] between consecutive frames:
@@ Appendix D More Quantitative Results
Aesthetic score follows the setup in VBench [huang2024vbench], using CLIP embeddings passed through a linear layer trained to predict human-rated aesthetic quality from a large-scale dataset of labeled images [schuhmann2022laion].
Dynamic degree quantifies how often videos exhibit meaningful motion. For each video, we compute optical flow using RAFT [teed2020raft] and identify the top 5 percent of pixels with the highest flow magnitude per frame. Frames exceeding a motion threshold are labeled as dynamic; a video is then classified as dynamic if a sufficient proportion of its frames meet this criterion. The dynamic degree of a group of videos is defined as the percentage of videos that are labeled dynamic. We show extra qualitative results in Table 10. Across these quantitative measures, all methods perform comparably, indicating that our method can successfully transfer reference effects without sacrificing overall video quality.
@@
Table 5: List of LoRA models and their corresponding text prompts used in this work. Each LoRA entry spans multiple lines for readability.

### v4

@@ 4.2 Baselines
Image-to-Video Baselines. We benchmark against the Wan 2.1 image-to-video model [74] and the Wan VACE model [36], which represent current state-of-the-art text-guided video generation systems.
Video-to-Video Baselines. For the video-to-video task, we evaluate two configurations based upon the Wan VACE [36] conditional model and one native text-based video editing model. Pose-conditioned Wan VACE: In this setup, we condition the Wan VACE on the first frame of the input video, intermediate poses extracted from the input video [81], and the corresponding text prompt. Depth-conditioned: In this setup, we condition the Wan VACE on the first frame of the input video, intermediate depth maps [60] from the input video, and the same text prompt. Lucy Edit [26]: Lucy Edit is a text-based video editing model built on Wan 2.2 [26], which directly takes an input video and a text instruction to produce an edited output.
@@ 4.3 Qualitative Results
Video-to-Video Effects. Qualitative comparisons for the video-to-video task are shown in Figure 7. Wan VACE variants either overfit to the input video or introduce artifacts, while Lucy Edit produces static, frame-invariant edits that lack temporal evolution. By conditioning on the full reference video, our method consistently reproduces complex time-varying transformations such as gradual color shifts or structural morphing while maintaining spatial consistency and motion alignment with the input.
@@
Figure 7: Results for our reference-based video-to-video setup. We compare three baselines: (1) Wan VACE + Pose: a conditional diffusion transformer [36] using the first input frame and extracted poses; (2) Wan VACE + Depth: using the first frame and extracted depths with Wan VACE; and (3) Lucy Edit [26], a video editing model fine-tuned from an image-to-video model. Without reference video information, all baselines fail to follow the reference from a prompt alone. Wan VACE methods overfit to the input or add irrelevant edits, while Lucy Edit applies uniform static edits across frames. Our method successfully follows the reference effect.
@@
User Study For a subjective task such as reference-based video editing or image-to-video generation, it is difficult to determine any single quantitative metric that fully captures success. We therefore conduct a comprehensive user study as the primary evaluation of our method across three tasks: neural video-to-video effect transfer (V2V), image-to-video effect transfer (I2V), and code-based temporal effects. Annotators on Amazon Mechanical Turk performed two-alternative forced choice (2AFC) pairwise comparisons between our method and several baselines, including Lucy Edit, VACE (Depth/Pose/I2V variants), Wan2.1, and an ablation that drops reference effect video conditioning (No Ref). Each comparison was evaluated along three dimensions: Reference Video Adherence (RVA), Input Video Adherence (IVA), and Overall Match (OM), where OM asks which output best applies the reference effect while preserving the input video. For I2V, we evaluate on RVA and OM. As shown in Table 1, our method achieves consistent preference over all baselines in V2V across all three dimensions, with win rates of 57.0–62.3% in RVA, 53.7–60.7% in IVA, and 60.0–67.2% in OM. Code-based effects show even stronger results, with win rates of 59.5–65.7% in RVA and 61.3–65.3% in OM. For I2V (Table 2), we observe win rates of 56.8–57.4% in RVA and 61.8–63.4% in OM against all baselines. All reported win rates exceed the 50% chance level, standard deviations confirm statistical reliability. The closest statistic in terms of preference is the input video adherence for our method without reference input video, which is expected as it uses the same model as our full method, and the test is simply adherence to the input video which both versions take as input. These results validate our approach’s effectiveness in effect transfer while preserving input content fidelity across multiple generation modalities. In total, we collected over 15000 annotations from over 500 unique annotators. For more details on the user study setup, please refer to Section C of the Appendix.
@@ 4.4 Quantitative Metrics
Conditioning Input Similarity. Since our validation sets do not include ground-truth triplets, we measure the similarity of our generated outputs to both the input and reference videos as a proxy for task success. To quantify this, we employ VideoPrism [87], a large-scale video embedding model pretrained on over 500M clips, to compute feature-space similarities between videos. For the reference + image-to-video case, we instead measure the average similarity between the generated video and its first-frame conditioning using CLIP [59] embeddings. Results are summarized in Tables 3 and 4. Our method consistently achieves higher similarity to the reference video than all baselines, confirming that it effectively incorporates temporal and stylistic information from the reference. Interestingly, baseline models often exhibit slightly higher similarity to the input, likely reflecting a form of under-editing. This phenomenon is also visible in Figure 7 (left) and Figure 6 (bottom), where baselines preserve input structure but fail to reproduce the desired temporal evolution.
@@ 4.5 Controllable Real Video Editing
When editing real videos, users may want to control the level to which the original video is maintained, as well as the amount that the reference effect video effects the final generation. To this end, we follow previous work [27, 8, 37, 42, 41, 71] and combine multiple classifier-free guidance directions during inference. Specifically, we consider applying guidance for text, input video, and reference effect video, normalizing latents as recomended in previous work [42]. This gives:
@@
Table 3: I2V Similarities. First Frame Sim is average clip [59] similarity to the input frame, and Ref Sim. is video embedding similarity [87] between the output video and the reference effect video
@@
Table 4: Neural and Code-Based V2V Similarities. Input Sim. is video embedding similarity [87] between the output video and the input video, and Ref Sim. is the video embedding similarity between the output video and the reference effect video.
@@ Appendix B Dataset Details
We source the Image-to-Video data from open-source LoRA effects trained on top of the Wan 2.1 Image-to-Video model [74], available on Hugging Face [1]. A comprehensive list of LoRA effects and their corresponding captions is provided in Table 5.
We use Qwen Image [78] and Qwen Image Edit [78], together with Wan 2.1 VACE-1.3B [74], to implement our text-guided video editing pair creation pipeline, as described in Figure 4 of the main text. We categorize our effect types and general effect categories in Table 6.
For the code-based effects, we source data from the grounding subset of the Senorita dataset [89]. Specifically, we filter for human-centric videos and include only grounding results with human-centric masks, yielding a collection of over 100K videos. The set of objects considered is listed in Table 7. In the original dataset, input videos are used to predict grounding masks; we repurpose these masks to obtain foreground or object and background segmentations, which we use to augment our effect pipelines. We manually curate code across a diverse set of effects and temporal transitions to sample from when creating triplets of the form (reference video, input video, output video). The full set of effect types and their corresponding hyperparameters is provided in Table 8, and the full set of temporal effect types and their corresponding hyperparameters is given in Table 9.
Appendix C User Study Details
We conduct our user study on Amazon Mechanical Turk (MTurk) using a two-alternative forced choice (2AFC) paradigm. In each trial, annotators are shown an input video and a reference effect video, followed by two output videos (one from our method and one from a baseline) presented side-by-side in randomized left/right order. Annotators select which output better satisfies the evaluation criterion. We evaluate three criteria in separate experiments: (1) Reference Video Adherence (RVA): which output better applies the reference effect; (2) Input Video Adherence (IVA): which output better preserves the input video; and (3) Overall Match (OM): which output best applies the reference effect while preserving the input video. Each experiment compares our method against a single baseline on a single evaluation dimension. Each experiment consists of 30 HITs (Human Intelligence Tasks), where each HIT is completed by a unique annotator. Within each HIT, annotators evaluate a subset of all available video pairs for that experiment, with the presentation order and left/right assignment randomized per HIT to mitigate ordering and side biases. To ensure annotation quality, each HIT begins with 5 practice trials with immediate feedback. In practice trials, annotators are shown a clearly correct and clearly incorrect output and must select the correct one; feedback is provided after each practice trial to calibrate annotators before they proceed to the main evaluation. We report win rates as the percentage of trials in which annotators preferred our method, where 50 percent represents chance performance. To estimate uncertainty, we perform bootstrap resampling with 10,000 iterations, resampling both annotators and trials with replacement. We report the standard deviation of the bootstrap distribution as our measure of statistical variability. Across all experiments, our method achieves win rates consistently above 50 percent, with bootstrap standard deviations typically in the range of 2–5 percent, confirming the statistical reliability of the observed preferences. See the setup in Figure 10.
Figure 10: User Study Example. Example of a task shown to users. Users are prompted with the input video and reference video, and asked to pick which of the videos shown are better with resepct to either input video adherence, reference video adherence, or overall match to the task.
Appendix D More Implementation Details
@@
Appendix E More Quantitative Results
We further assess video fidelity and perceptual quality using established metrics from VBench [34], including motion smoothness [74], aesthetic score [45], and dynamic degree [72, 34].
Motion quality measures temporal consistency by computing the average cosine similarity [59] between consecutive frames:
@@ Appendix E More Quantitative Results
Aesthetic score follows the setup in VBench [34], using CLIP embeddings passed through a linear layer trained to predict human-rated aesthetic quality from a large-scale dataset of labeled images [66].
Dynamic degree quantifies how often videos exhibit meaningful motion. For each video, we compute optical flow using RAFT [72] and identify the top 5 percent of pixels with the highest flow magnitude per frame. Frames exceeding a motion threshold are labeled as dynamic; a video is then classified as dynamic if a sufficient proportion of its frames meet this criterion. The dynamic degree of a group of videos is defined as the percentage of videos that are labeled dynamic. We show extra qualitative results in Table 10. Across these quantitative measures, all methods perform comparably, indicating that our method can successfully transfer reference effects without sacrificing overall video quality.
@@
Table 5: Subset of list of LoRA models and their corresponding text prompts used in this work. Each LoRA entry spans multiple lines for readability.